What quantity of 80 per cent acid solution must be mixed with a 30 per cent solution to produce 1050 mL of a 50 per cent solution?
mL of acid solution =

Answer by mananth(16946)

percent ---------------- quantity
Acid I 80 ---------------- x
Acid II 30 ---------------- 1050 - x
Total 50.00% ---------------- 1050

80x+30(1050-x)= 50*1050
80x+31500-30x=52500
80x-30x =52500-31500
50x=21000
/50
x=420 ml 80.00%
630 ml 30.00%